# Twisted Superfields

### v8

@@ 4 Unification and Anomalies
where the $m=2$ flavor 2,3 singlet scalars have zero hypercharge. Between $\tilde{M}_{\nu}$ and the electroweak scale, the 6 neutrinos are reduced to 3, the 6 charged lepton flavors are reduced to 4, and the 12 non-singlet scalars are reduced to 11. So between $\tilde{M}_{\nu}$ and $M_{Z}$, $b_{Y}=403/60$.
Running the $Y$ coupling up to $\tilde{M}_{\nu}$ and then up to $\Lambda_{U}$, the hypercharge coupling at the unification scale is
@@
So in the simplified scenario presented here, the three couplings come close to unifying at the same scale. Different assumptions about masses of the heavy charged leptons could lead to $\alpha_{Y}^{-1}\left(\Lambda_{U}\right)\simeq 28$. In that case, $g_{+}$ and $g$ would take the same value at the unification scale (motivating that assumption in section 2).
As mentioned above, $g_{Y^{\prime}}$ also runs more quickly below the scale of $m_{t}$ due to a combination of a larger beta function and the top-quark-induced effective anomaly. It is proposed that if the effect of the effective anomaly was fully taken into account, a $g_{Y^{\prime}}$ that unified with the other couplings at the unification scale could potentially run down fast enough (transforming to $g_{Z^{\prime}}$ at $M_{Z}$) in order to reach the extremely small values required to reproduce the measurements of $Z^{\prime}$ boson interactions described in section 5. Future work would be required to determine whether the suggested scenario can be supported by detailed calculations.
@@ 5 Experimental Implications
The model proposed in this paper is very different than the Standard Model. To truly define this model, more detailed calculations would need to be performed that are outside the scope of this paper. But even in the absence of these calculations, a number of statements can be made about features (e.g. masses, couplings) the model would have to have in order to reproduce experimental data.

### v9

@@ 4 Unification and Anomalies
where the $m=2$ flavor 2 singlet scalars have zero hypercharge. Between $M_{\nu}^{(2)}$ and the electroweak scale, the 6 lepton families are reduced to 5, the leptoquarks are excluded, and one $m=2$ scalar doublet and singlet is also excluded. So between $M_{\nu}^{(2)}$ and $M_{Z}$, $b_{Y}=131/30$.
Running the $Y$ coupling up to $M_{\nu}^{(2)}$ and then up to $\Lambda_{U}$, the hypercharge coupling at the unification scale is
@@
So in the simplified scenario presented here, the three couplings come close to unifying at the same scale. If they do indeed unify, then $g_{+}$ and $g$ take the same value at the unification scale (motivating that simplifying assumption in section 2).
If a different scale is taken for $M_{\nu}^{(2)}$ and different assumptions are made about various masses (e.g. leptoquarks), there are scenarios where $g_{+}(\Lambda_{U})$ does not have the same value as $g_{2}(\Lambda_{U})=g_{3}(\Lambda_{U})$. That makes the expressions for $Y^{\prime}$ and $Z^{\prime}$ group structure in sections 2 and 5 more complicated, but it does not qualitatively change the arguments made.
@@ 5 Experimental Implications
The model proposed in this paper is significantly different than the Standard Model. To definitively test this model, more detailed calculations would need to be performed that are outside the scope of this paper. But even in the absence of these calculations, a number of statements can be made about features (e.g. masses, couplings) the model would have to have in order to reproduce experimental data.